The area of the rectangle is 68 square meters. What are the width and length of the rectangle ? Give answers both exactly and approximately(to the nearest tenth)
In this problem all it shows is a rectangle with an "x" to the left side and "4x" on the bottom. not sure at all what to do.

Thanks for the help. Have a great day! :)

Answer by Boreal(15235)

A width and a length multiplied together will give you the area.
x*4x=4x^2, and that equals 68 m^2
divide by 4 and x^2=17 m^2
x=sqrt(17)=4.1 m
4x=16.4 m
Those are the dimensions, width by convention being the shorter and length being the longer.
4.1*16.4=67.24 m^2, a little smaller because we rounded down.